Solve the quadratic equation by factoring and applying the property "ab=0 if and only if a=0 or b=0"
(x^2)-4x-21=0

Answer by fractalier(6550)

First we factor the trinomial
x^2 - 4x - 21 = 0
into its binomial factors:
(x - 7)(x + 3) = 0
Now if two quantities multiply together to make zero, one (or both) has to be zero...
Thus x - 7 = 0 and x + 3 = 0 which leads to
x = 7 or x = -3